OpenStudy (anonymous):

I said $46.50 and it was wrong. Help please??? Samantha McQuade bought a blender and gave the store's clerk a $50 bill and a $1 bill. She received $.72 in change. If the sales tax rate in her community is 7.5%, what was the selling price of the blender before taxes?

jimthompson5910 (jim_thompson5910):

How much is the blender after taxes?

OpenStudy (anonymous):

$50.28 ?

jimthompson5910 (jim_thompson5910):

Now let x be the price before taxes (Original price before taxes) + (7.5% of (Original price before taxes)) = Final Price x + 7.5% of x = 50.28 x + 0.075x = 50.28 1x + 0.075x = 50.28 1.075x = 50.28 x = ???

OpenStudy (anonymous):

so 46.77 ??

OpenStudy (larseighner):

107.5 % of the sale price = $51 - $0.72 1.075 x = 50.28 x = 50.28/1.075 x = 46.77 (approx.)

jimthompson5910 (jim_thompson5910):

correct, $46.77

OpenStudy (anonymous):

thank you both